# Navigating Distribution Shifts in Medical Image Analysis: A Survey

### v1

@@ VI-B2 Model Design
Batch Normalization. Batch Normalization has been widely explored in adaptation tasks as normalization statistics are associated with the domain distribution. They can be directly obtained through pre-trained model and taken as the source information. Given a mini-batch =$\mathcal{B}=\{x_{n}\}_{n=1}^{N}$ where $x_{n}\in\mathbb{R}^{F}$ is a feature vector (with $F$ denoting the number of feature channels and $N$ the batch size), BN normalizes each feature dimension $f$ as follows:
@@
$\hat{x}_{n,f}=\frac{x_{n,f}-\mu_{\mathcal{B},f}}{\sqrt{\sigma_{\mathcal{B},f}+\epsilon}}\cdot\gamma_{f}+\beta_{f}$ (12)
@@
where $\mu_{\mathcal{B},f}$ and $\sigma_{\mathcal{B},f}$ are the running mean and variance for the $f$-th feature of mini-batch $\mathcal{B}$, respectively. The parameters $\gamma_{f}$ and $\beta_{f}$ are the learned scale and shift factors for affine transformation, with $\epsilon$ being a small-offset to avoid division by zero. [200] proposes an exponential decay scheme for the normalization statistics in adaptation stage to gradually learn the target domain-specific mean and variance. [199] aligns the source and target normalization statistic discrepancy for learning a prompt to make the target inputs be treated as the source. More recently, [201] explores domain-specific and shareable batch normalization statistics for adaptive BN-based adaptation, while [202] proposes to incorporate the concept of class diversity to address more realistic mini-batch problem.
@@
Novel Strategies and Structures. The field has seen several structural innovations aimed at overcoming specific adaptation challenges. [203] introduces an auxiliary rotation classifier to improve adaptation via self-training. Similarly, [204] utilizes multiple diverse classifiers to address test label distribution shifts, and [205] employs decoder duplication during the adaptation stage to ensemble diverse target inputs. Y-shaped architectures with dual decoders are used for enhanced denoising and segmentation [206, 207].  [208] further develops a supplementary network to adaptively combined with the main outputs during inference.
VI-B3 Optimization Strategy
@@
Entropy Minimization. Entropy minimization is widely-used to handle unlabeled data. Mathematically, the entropy of a prediction can be expressed as follows:
$H(p)=-\sum_{i=1}^{C}p_{i}\log p_{i}$ (13)
@@
where $p$ represents the predicted probability distribution over $C$ classes, and $p_{i}$ is the probability of the $i$-th class predicted by the model. The goal is to minimize this entropy $H(p)$ across the dataset, thereby encouraging the model to produce more decisive outputs. This approach is first introduced by Tent [184] into general TTA tasks, which proposes minimizing the mean entropy over the test batch to update the affine parameters of the batch normalization layers in the pre-trained model. This strategy has been adopted in many cases [209, 200, 208, 210, 211] in medical TTAs.
@@
Dynamic Adjustment of Learning Rates. Dynamic adjustment of learning rates based on distribution shifts helps models adapt more effectively during test-training stages. For example, [27] proposes that samples with larger distribution shift should result in larger update. It makes adjustment by calculating the divergence between the model outputs and its nearest neighbors in a memory bank. [212] refines this strategy by assessing category-wise discrepancies with an uncertainty estimation module.
@@
Anatomical Information. Leveraging the anatomical information as a prior for loss design offers a promising direction for enhancing the accuracy and reliability. For instance, [213] utilizes a shape dictionary, integrating general semantic shapes extracted from source data. [214] incorporates the shape information with the signed distance field which measures the distance between any pixel to the nearest object boundary and the relative position. [210, 211] leverages the class-ratio as a supervision which is estimated from anatomical knowledge available in the clinical literature. [215] uses anatomically-derived loss functions that penalizes unrealistic bone lengths and joint angles in 3D pose estimation. [216] proposes contour regularization loss for constraining the continuity and connectivity. [217] expands the lesion click (i.e., the center of the nodule) into an ellipsoid mask, and use it as the supervised information for test training.
VII Domain Generalization
@@
Domain Generalization (DG) is an advanced deep learning technique designed to prepare models for handling unseen, out-of-distribution data. This challenge is especially relevant in MedIA, where real-world operational constraints often make target datasets from new domains inaccessible or unknown. In such cases, DG methods become essential, enabling models to generalize from available source data to new environments without prior exposure to specific target data. This proactive approach ensures that medical models remain robust and accurate, ready to cope with potential unfamiliar environments. Furthermore, unlike other categories that are typically divided from simple to complex based on the label availability of target data, DG assumes that no target data is available. The complexity of tasks within this category primarily hinges on the nature of the source data. Specifically, DG techniques can be divided into two main types: Multi-source Domain Generalization (MDG) and Single-source Domain Generalization (SDG). MDG capitalizes on the diversity of multiple source datasets to extract and decouple domain-invariant and domain-specific features, thereby enhancing the model’s generalizability using the domain-invariant component. Conversely, SDG, limited to a single source, faces greater challenges and often relies on additional data augmentation strategies to increase the model’s generalization capabilities under more restrictive conditions.
VII-A Multi-source Domain Generalization
@@
VII-A1 Model Design
Meta-learning. Meta-learning [218] is a powerful strategy for enhancing model generalization across unknown data distributions. This approach involves simulating domain shifts during training through “episodes”, where data from multiple sources is split into meta-train $\mathcal{D}_{\text{train}}$ and meta-test $\mathcal{D}_{\text{test}}$ sets. This split mirrors real-world domain shifts, preparing the model for new domains or distributions. The model first learns from the meta-train set and is then tested on the meta-test set to evaluate its adaptability to new situations. Adjustments are made based on its performance to enhance its generalization capabilities. This process is formulated as:
$\small\phi^{*}=\text{{MetaLearn}}(\mathcal{D}_{\text{train}}),\theta^{*}=\text{{Learn}}(\mathcal{D}_{\text{test}};\phi^{*})$ (14)
where $\phi^{*}$ denotes the meta-learned parameters, which are then used to learn the task-specific model parameters $\theta^{*}$ on the meta-test set. Following this framework, [15] introduces a shape-aware meta-learning scheme that incorporates anatomical integrity, [28] combines meta-learning with style-feature flow generation for confounding factors elimination, and [219] uses style-transferred images as meta-tests, designing a new boundary-oriented objective for meta-optimization considering the specific challenges in medical image segmentation.
VII-A2 Optimization Strategy
Shape-based Regularization. Shape-based Regularization is a powerful tool, harnessing the continuous and coherent nature of anatomical structures and the domain-invariant characteristics of their contours. Except for combined with the meta-learning approaches [15, 28, 219] for supervision in meta-test optimization, some methods directly use the anatomical knowledge as prior information during training. For example, [220] integrates fixed Sobel kernels for contour enhancement and a convolutional autoencoder for learning anatomical priors, which inversely projects the mask and prediction to the feature space for further alignment.
Latent Space Regularization. Latent Space Regularization focuses on modeling inter-domain relationships and perform regularization in the latent feature space to promote generalization. Notably, [221] introduces a rank regularization term to constrain the complexity of feature representations and restrict the latent features to follow a pre-defined prior distribution, while [222] implements semantic feature regularization during the meta-test phase with dual losses that maintain global inter-class relationships and tighten intra-class features.
VII-B Single-source Domain Generalization
@@
VII-B1 Data Management
Pixel-level Augmentation. Pixel-level Augmentation techniques directly manipulate the pixel values. This method is primarily based on the premise that variations in imaging modalities, acquisition protocols, and hardware can induce significant discrepancies in image characteristics such as texture, intensity, and contrast. For example, [223] introduced BigAug, a deep stacked general transformation approach to systematically evaluate augmentation effects on model generalization. Specialized approaches within the medical field, such as the use of Bézier Curves by [29] to address gray-scale discrepancies, and the simulation of MRI distortions by [224], focus on medical-specific image traits. [225] uses causal inference methods to reflect acquisition shifts and [226] explores category-level augmentation based on class-level representation invariance. [227] combines the augmentation strategies both in [223] and [226]. [228] expands the style space through adversarial training and finds the worst-case style composition to generate the samples. [229] further refines this strategy by introducing randomness to the generated domain through a adaptive instance normalization block, so that the changes are limited to the textures.
Feature-level Augmentation. Some augmentation strategies delves deeper into the model’s internal workings, focusing on the manipulation of learned feature representations. For instance, [230] masks features channel-wisely and spatially to generate diverse challenging samples.
VIII Future Directions
@@
VIII-A Continuous Learning Systems for MedIA
@@
In medical imaging analysis, continuous learning systems [231, 232] are essential for adapting to dynamic environments where data distributions evolve over time. As medical practices advance and new clinical data becomes available, models must accommodate these shifts without the need for complete retraining. For instance, data collected over different time periods – such as imaging acquired from new devices, updated imaging protocols, or the emergence of previously unseen disease variants – leads to temporal distribution shifts that challenge conventional static models. Continuous learning addresses these shifts by incrementally updating the model as new data arrives, ensuring that it remains relevant to evolving clinical scenarios. Importantly, these systems are designed to combat catastrophic forgetting, where a model loses performance on previously learned tasks as it incorporates new knowledge. Techniques such as rehearsal, regularization, and memory-based strategies allow continuous learning models to maintain stable performance across a broad spectrum of conditions, ultimately improving adaptability in clinical workflows where the nature of data is in constant flux.
VIII-B Harnessing Vision-Foundation Models for MedIA
@@
The utilization of Vision-Foundation Models (VFMs) presents a promising direction for mitigating distribution shifts in MedIA. For instance, existing research has led to the development of MedSAM [10] and MedCLIP [233], which are fine-tuned versions of SAM [234] and CLIP [235], respectively. MedSAM excels in universal segmentation, allowing it to adapt to diverse medical datasets and imaging protocols, thereby showing its robustness against variations in data distribution. MedCLIP leverages zero-shot learning capabilities, aiming to recognize and analyze unseen medical images based on the embeddings of created prompts for each disease class, which is invaluable in clinical settings with limited labeled data. Further investigation could be the combination of these models for innovative solutions, such as a comprehensive diagnostic tool that integrates segmentation and contextual analysis, providing real-time insights that adjust to new imaging conditions and patient histories. Moreover, leveraging existing large models to enhance the functionality of another model, thereby reducing the impact of distribution shifts, is a promising direction worth exploring. For example, utilizing CLIP to generate visually descriptive sentences related to the segmentation target could enable SAM to effectively perform zero-shot medical segmentation [236]. Future research should explore these synergies, potentially leading to multi-task learning frameworks that simultaneously address various medical tasks, thus enhancing diagnostic accuracy and personalizing treatment strategies. By pursuing these avenues, we can unlock the full potential of VFMs to create resilient and adaptable systems that significantly improve healthcare outcomes.
VIII-C Multi-task/modal Learning under Distribution Shifts
@@
Multi-task [237] and multi-modal [238] learning frameworks offer a promising avenue for addressing the complexities inherent in medical image analysis, particularly when data arises from diverse sources or modalities. These models, designed to jointly process data such as medical images, textual reports, and real-time procedural videos, facilitate a more comprehensive understanding of patient conditions. For example, a multi-modal system could simultaneously perform tumor segmentation from MRI, extract pertinent clinical information from patient records, and analyze surgical video footage to assess tissue responses. Such integrative models are robust against distributional shifts that occur between different data types, ensuring that the system can handle variable quality and types of inputs. Moreover, by leveraging shared representations across tasks and modalities, the model is better equipped to capture general, robust features rather than task-specific nuances, which might overfit to a single type of data.
IX Conclusion

### v2

@@ 7.2.2. Model Design
Batch Normalization has been widely explored in adaptation tasks as normalization statistics are associated with the domain distribution. They can be directly obtained through pre-trained model and taken as the source information. Given a mini-batch =$\mathcal{B}=\{x_{n}\}_{n=1}^{N}$ where $x_{n}\in\mathbb{R}^{F}$ is a feature vector (with $F$ denoting the number of feature channels and $N$ the batch size), BN normalizes each feature dimension $f$ as follows:
@@
(9) $\hat{x}_{n,f}=\frac{x_{n,f}-\mu_{\mathcal{B},f}}{\sqrt{\sigma_{\mathcal{B},f}+\epsilon}}\cdot\gamma_{f}+\beta_{f}$
@@
where $\mu_{\mathcal{B},f}$ and $\sigma_{\mathcal{B},f}$ are the running mean and variance for the $f$-th feature of mini-batch $\mathcal{B}$, respectively. The parameters $\gamma_{f}$ and $\beta_{f}$ are the learned scale and shift factors for affine transformation, with $\epsilon$ being a small-offset to avoid division by zero. (Liu et al., 2021g) proposes an exponential decay scheme for the normalization statistics in adaptation stage to gradually learn the target domain-specific mean and variance. (Hu et al., 2022c) aligns the source and target normalization statistic discrepancy for learning a prompt to make the target inputs be treated as the source. More recently, (Liu et al., 2023c) explores domain-specific and shareable batch normalization statistics for adaptive BN-based adaptation, while (Su et al., 2024) proposes to incorporate the concept of class diversity to address more realistic mini-batch problem.
7.2.2.2 Novel Strategies and Structures.
@@
The field has seen several structural innovations aimed at overcoming specific adaptation challenges.  (Feng et al., 2023b) introduces an auxiliary rotation classifier to improve adaptation via self-training. Similarly, (Ma et al., 2022) utilizes multiple diverse classifiers to address test label distribution shifts, and (Wu et al., 2023a) employs decoder duplication during the adaptation stage to ensemble diverse target inputs. Y-shaped architectures with dual decoders are used for enhanced denoising and segmentation (Wen et al., 2024; Karani et al., 2021).  (Zhang et al., 2024) further develops a supplementary network to adaptively combined with the main outputs during inference.
7.2.3. Optimization Strategy
7.2.3.1 Entropy Minimization.
@@ 7.2.3. Optimization Strategy
Entropy minimization is widely-used to handle unlabeled data. Mathematically, the entropy of a prediction can be expressed as follows:
(10) $H(p)=-\sum_{i=1}^{C}p_{i}\log p_{i}$
@@
where $p$ represents the predicted probability distribution over $C$ classes, and $p_{i}$ is the probability of the $i$-th class predicted by the model. The goal is to minimize this entropy $H(p)$ across the dataset, thereby encouraging the model to produce more decisive outputs. This approach is first introduced by Tent (Wang et al., 2021) into general TTA tasks, which proposes minimizing the mean entropy over the test batch to update the affine parameters of the batch normalization layers in the pre-trained model. This strategy has been adopted in many cases (Hong et al., 2022b; Liu et al., 2021g; Zhang et al., 2024; Bateson et al., 2022a, b) in medical TTAs.
7.2.3.2 Dynamic Adjustment of Learning Rates.
@@
Dynamic adjustment of learning rates based on distribution shifts helps models adapt more effectively during test-training stages. For example, (Yang et al., 2022a) proposes that samples with larger distribution shift should result in larger update. It makes adjustment by calculating the divergence between the model outputs and its nearest neighbors in a memory bank. (Zhang et al., 2023b) refines this strategy by assessing category-wise discrepancies with an uncertainty estimation module.
7.2.3.3 Anatomical Information.
@@
Leveraging the anatomical information as a prior for loss design offers a promising direction for enhancing the accuracy and reliability. For instance, (Liu et al., 2022a) utilizes a shape dictionary, integrating general semantic shapes extracted from source data.  (Zhu et al., 2023a) incorporates the shape information with the signed distance field which measures the distance between any pixel to the nearest object boundary and the relative position. (Bateson et al., 2022a, b) leverages the class-ratio as a supervision which is estimated from anatomical knowledge available in the clinical literature. (Bigalke et al., 2023) uses anatomically-derived loss functions that penalizes unrealistic bone lengths and joint angles in 3D pose estimation. (Hu et al., 2021) proposes contour regularization loss for constraining the continuity and connectivity. (Li et al., 2023c) expands the lesion click (i.e., the center of the nodule) into an ellipsoid mask, and use it as the supervised information for test training.
Figure 6. Domain Generalization in MedIA: It prepares models for unseen data by generalizing from source datasets without target data access. This proactive approach ensures robustness in unfamiliar environments. The shown flowchart is Single-source Domain Generalization.
8. Domain Generalization
@@
Domain Generalization (DG) is an advanced deep learning technique designed to prepare models for handling unseen, out-of-distribution data. This challenge is especially relevant in MedIA, where real-world operational constraints often make target datasets from new domains inaccessible or unknown. In such cases, DG methods become essential, enabling models to generalize from available source data to new environments without prior exposure to specific target data (see toy example in Fig. 6). This proactive approach ensures that medical models remain robust and accurate, ready to cope with potential unfamiliar environments. Furthermore, unlike other categories that are typically divided from simple to complex based on the label availability of target data, DG assumes that no target data is available. The complexity of tasks within this category primarily hinges on the nature of the source data. Specifically, DG techniques can be divided into two main types: Multi-source Domain Generalization (MDG) and Single-source Domain Generalization (SDG). MDG capitalizes on the diversity of multiple source datasets to extract and decouple domain-invariant and domain-specific features, thereby enhancing the model’s generalizability using the domain-invariant component. Conversely, SDG, limited to a single source, faces greater challenges and often relies on additional data augmentation strategies to increase the model’s generalization capabilities under more restrictive conditions.
8.1. Multi-source Domain Generalization
@@
8.1.1. Model Design
8.1.1.1 Meta-learning.
Meta-learning (Hospedales et al., 2021) is a powerful strategy for enhancing model generalization across unknown data distributions. This approach involves simulating domain shifts during training through “episodes”, where data from multiple sources is split into meta-train $\mathcal{D}_{\text{train}}$ and meta-test $\mathcal{D}_{\text{test}}$ sets. This split mirrors real-world domain shifts, preparing the model for new domains or distributions. The model first learns from the meta-train set and is then tested on the meta-test set to evaluate its adaptability to new situations. Adjustments are made based on its performance to enhance its generalization capabilities. This process is formulated as:
(11) $\small\phi^{*}=\text{{MetaLearn}}(\mathcal{D}_{\text{train}}),\theta^{*}=\text{{Learn}}(\mathcal{D}_{\text{test}};\phi^{*})$
where $\phi^{*}$ denotes the meta-learned parameters, which are then used to learn the task-specific model parameters $\theta^{*}$ on the meta-test set. Following this framework, (Liu et al., 2020a) introduces a shape-aware meta-learning scheme that incorporates anatomical integrity, (Li et al., 2022a) combines meta-learning with style-feature flow generation for confounding factors elimination, and (Liu et al., 2021a) uses style-transferred images as meta-tests, designing a new boundary-oriented objective for meta-optimization considering the specific challenges in medical image segmentation.
8.1.2. Optimization Strategy
8.1.2.1 Shape-based Regularization.
Shape-based Regularization is a powerful tool, harnessing the continuous and coherent nature of anatomical structures and the domain-invariant characteristics of their contours. Except for combined with the meta-learning approaches (Liu et al., 2020a; Li et al., 2022a; Liu et al., 2021a) for supervision in meta-test optimization, some methods directly use the anatomical knowledge as prior information during training. For example, (Pham et al., 2020) integrates fixed Sobel kernels for contour enhancement and a convolutional autoencoder for learning anatomical priors, which inversely projects the mask and prediction to the feature space for further alignment.
8.1.2.2 Latent Space Regularization.
Latent Space Regularization focuses on modeling inter-domain relationships and perform regularization in the latent feature space to promote generalization. Notably, (Li et al., 2020c) introduces a rank regularization term to constrain the complexity of feature representations and restrict the latent features to follow a pre-defined prior distribution, while (Dou et al., 2019a) implements semantic feature regularization during the meta-test phase with dual losses that maintain global inter-class relationships and tighten intra-class features.
8.2. Single-source Domain Generalization
@@
8.2.1. Data Management
8.2.1.1 Pixel-level Augmentation.
Pixel-level Augmentation techniques directly manipulate the pixel values. This method is primarily based on the premise that variations in imaging modalities, acquisition protocols, and hardware can induce significant discrepancies in image characteristics such as texture, intensity, and contrast. For example, (Zhang et al., 2020b) introduced BigAug, a deep stacked general transformation approach to systematically evaluate augmentation effects on model generalization. Specialized approaches within the medical field, such as the use of Bézier Curves by (Zhou et al., 2022a) to address gray-scale discrepancies, and the simulation of MRI distortions by (Kamraoui et al., 2022), focus on medical-specific image traits. (Ouyang et al., 2022) uses causal inference methods to reflect acquisition shifts and (Su et al., 2023a) explores category-level augmentation based on class-level representation invariance. (Hu et al., 2023) combines the augmentation strategies both in (Zhang et al., 2020b) and (Su et al., 2023a). (Chen et al., 2022b) expands the style space through adversarial training and finds the worst-case style composition to generate the samples. (Xu et al., 2022c) further refines this strategy by introducing randomness to the generated domain through a adaptive instance normalization block, so that the changes are limited to the textures.
8.2.1.2 Feature-level Augmentation.
Some augmentation strategies delves deeper into the model’s internal workings, focusing on the manipulation of learned feature representations. For instance, (Chen et al., 2021a) masks features channel-wisely and spatially to generate diverse challenging samples.
9. Future Directions
@@
9.1. Continuous Learning Systems for MedIA
@@
In medical imaging analysis, continuous learning systems (Zhang et al., 2023c; González et al., 2023) are essential for adapting to dynamic environments where data distributions evolve over time. As medical practices advance and new clinical data becomes available, models must accommodate these shifts without the need for complete retraining. For instance, data collected over different time periods – such as imaging acquired from new devices, updated imaging protocols, or the emergence of previously unseen disease variants – leads to temporal distribution shifts that challenge conventional static models. Continuous learning addresses these shifts by incrementally updating the model as new data arrives, ensuring that it remains relevant to evolving clinical scenarios. Importantly, these systems are designed to combat catastrophic forgetting, where a model loses performance on previously learned tasks as it incorporates new knowledge. Techniques such as rehearsal, regularization, and memory-based strategies allow continuous learning models to maintain stable performance across a broad spectrum of conditions, ultimately improving adaptability in clinical workflows where the nature of data is in constant flux.
9.2. Harnessing Vision-Foundation Models for MedIA
@@
The utilization of Vision-Foundation Models (VFMs) presents a promising direction for mitigating distribution shifts in MedIA. For instance, existing research has led to the development of MedSAM (Ma et al., 2024) and MedCLIP (Wang et al., 2022b), which are fine-tuned versions of SAM (Kirillov et al., 2023) and CLIP (Radford et al., 2021), respectively. MedSAM excels in universal segmentation, allowing it to adapt to diverse medical datasets and imaging protocols, thereby showing its robustness against variations in data distribution. MedCLIP leverages zero-shot learning capabilities, aiming to recognize and analyze unseen medical images based on the embeddings of created prompts for each disease class, which is invaluable in clinical settings with limited labeled data. Even when VFMs are not used for direct prediction, they can provide additional information to local models to assist in prediction and analysis. For example, VFMs can highlight regions of interest, enhancing the explainability and credibility of the models (Patr´ıcio et al., 2023; Hossain et al., 2023; Messina et al., 2022). Further investigation could be the combination of these models for innovative solutions, such as a comprehensive diagnostic tool that integrates segmentation and contextual analysis, providing real-time insights that adjust to new imaging conditions and patient histories. Moreover, leveraging existing large models to enhance the functionality of another model, thereby reducing the impact of distribution shifts, is a promising direction worth exploring. For example, utilizing CLIP to generate visually descriptive sentences related to the segmentation target could enable SAM to effectively perform zero-shot medical segmentation (Aleem et al., 2024). Future research should explore these synergies, potentially leading to multi-task learning frameworks that simultaneously address various medical tasks, thus enhancing diagnostic accuracy and personalizing treatment strategies. By pursuing these avenues, we can unlock the full potential of VFMs to create resilient and adaptable systems that significantly improve healthcare outcomes.
9.3. Multi-task/modal Learning under Distribution Shifts
@@
Multi-task (Zhao et al., 2023) and multi-modal (Cao et al., 2023) learning frameworks offer a promising avenue for addressing the complexities inherent in medical image analysis, particularly when data arises from diverse sources or modalities. These models, designed to jointly process data such as medical images, textual reports, and real-time procedural videos, facilitate a more comprehensive understanding of patient conditions. For example, a multi-modal system could simultaneously perform tumor segmentation from MRI, extract pertinent clinical information from patient records, and analyze surgical video footage to assess tissue responses. Such integrative models are robust against distributional shifts that occur between different data types, ensuring that the system can handle variable quality and types of inputs. Moreover, by leveraging shared representations across tasks and modalities, the model is better equipped to capture general, robust features rather than task-specific nuances, which might overfit to a single type of data.
10. Conclusion